OpenStudy (anonymous):

How would you find the hypotenuse r when given the opposite side whose value is 20 and the angle 45 degrees?

OpenStudy (anonymous):



OpenStudy (jdoe0001):

bear in mind the identity of \(\large tan(\theta) = \cfrac{opposite}{hypotenuse} \implies hypotenuse = \cfrac{opposite}{tan(\theta)}\)

OpenStudy (anonymous):

So is the answer 20?

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

tan doesn't equal opp/hyp

OpenStudy (anonymous):

yeah that's what i thought o.o I thought tan= opp/adj

OpenStudy (anonymous):

yea it does :)

OpenStudy (anonymous):

\[\sin(45)=\frac{ opp }{ hyp }\]

OpenStudy (anonymous):

in the end it was actually 20root2

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ 2 }=\frac{ 20 }{ hyp}\]

OpenStudy (anonymous):

Can you do it from there?

OpenStudy (whpalmer4):

you could do this all with Pythagorean theorem, no trig required. If 1 angle is 45, and another is 90, the 3rd angle must be 180-45-90 = 45. That implies that the two legs are both 20, so by PT: \[20^2+20^2=r^2\]\[r^2=2*20^2\]\[r=20\sqrt{2}\]

OpenStudy (anonymous):

As my dad would say, S.S.D.P.

OpenStudy (anonymous):

thanks for your help, I'm not liking trig >:T

OpenStudy (anonymous):

I did really well in trig, when I'm on I'll anyway I can. :)